OpenStudy (anonymous):

The population of a certain town increased by 2% from 35,000. What was the population after the increase?

OpenStudy (anonymous):

\[P' = P*(1+\frac{ r }{ 100 })^n\] where P' = population after n years P = present population

OpenStudy (anonymous):

P*r + P of course

OpenStudy (anonymous):

took me a second to figure out what the 100 was doing there :)

OpenStudy (anonymous):

might want to round up to avoid a fraction of a person walking about ....